# Optimizing Parameters of the LinDistFlow Power Flow Approximation for Distribution Systems

### v3

@@ V-A Algorithm Training
As in [19], we use $20$ power injection scenarios during our algorithm’s training phase. This selection is supported by a sensitivity analysis (see Fig. 3), which demonstrated diminishing returns in accuracy improvement with a larger number of scenarios, yielding a balance between model fidelity and training efficiency. These scenarios scale the nominal power injections at each bus by normally distributed multipliers with mean of one and standard deviation of $35\%$. For the nonlinear DistFlow solutions, we use PowerModels.jl [49] and OpenDSS with the OpenDSSDirect.py package for single- and three-phase networks, respectively, on a PACE computing node at Georgia Tech with a 24-core CPU and 32 GB of RAM. The training algorithm is implemented in Python 3 in a Jupyter Notebook using the TNC method from scipy.optimize.minimize with objective function (5), Jacobian $\mathbf{g}=[{\mathbf{g}^{D_{r}}}^{\top},{\mathbf{g}^{D_{x}}}^{\top},{\mathbf{g}^{\gamma}}^{\top},{\mathbf{g}^{\rho}}^{\top},{\mathbf{g}^{\varrho}}^{\top}]^{\top}$, convergence tolerance of $1\times 10^{-6}$ per unit, and iteration limit of $100$. [^1] Code is available at https://github.com/BabakTaheri1/OLDF
@@ V-B Performance Metrics
[…]
$\varepsilon_{\text{max}}^{[\text{model}]}$ $=\|\mathbf{v}^{[\text{model}]}-\mathbf{v}^{DF}\|_{\infty}$ (13)
$\varepsilon_{\text{avg}}^{[\text{model}]}$ $=\frac{1}{|\mathcal{M}||\mathcal{N}^{\prime}|}\|\mathbf{v}^{[\text{model}]}-\mathbf{v}^{DF}\|_{1}$ (14)
@@
Figure 7: Voltage profiles for the base case using the nonlinear DistFlow (OpenDSS in black), LinDistFlow with initial parameters (LDF in green), and LinDistFlow with optimized parameters (OLDF in red).
@@ V-D2 High Load Evaluation
Following the methodology described in [19], we generated high-load scenarios by scaling the base loads with a factor that ranges from $[-2,-1]\cup[1,2]$ at a granularity of $\frac{1}{14}$, yielding $30$ distinct test scenarios. As shown in Table II, our proposed OLDF algorithm surpasses the others in reducing the average voltage estimation error ($\varepsilon_{avg}$) across nearly all scenarios with the exception of the 22-bus case. For this case, our OLDF results were better than all except the PLPF approximation where the average error was still quite close ($0.00020$ p.u. for PLPF vs. $0.00028$ p.u. for OLDF). Regarding the maximum errors in the high-load scenarios, no individual approximation consistently dominated the others across all test cases. However, we note that summing the maximum errors across all test cases reveals that the OLDF parameters lead to the best performance in aggregate for this metric. These results show that OLDF parameters trained with scenarios around base-load conditions nevertheless perform well for high-load conditions. Furthermore, as an adaptive power flow, the OLDF approximation has the ability to tailor the parameters to perform even better for these conditions by including more training scenarios associated with high loading.
@@ V-E Computational Efficiency
The computation times for online uses of the optimized LinDistFlow parameters depend on the particular application for which they are employed. However, since the only changes are to the parameter values and not the mathematical form of the LinDistFlow expressions, online computation times with our optimized parameters should be comparable to existing LinDistFlow approximations. To show this, the average calculation times for the $10,000$ random load scenarios in the experiment in Section V-D3 ranged from $0.0004$ to $0.1156$ seconds, with differences of less than $2$% between the LinDistFlow using traditional parameters and our optimized parameters.
@@ V-F Topology Analysis
Addressing topology changes with the LinDistFlow approximation can be approached in various ways. For example, the traditional LinDistFlow approximation allows for straightforward topology adjustments by updating the $\mathbf{D}_{r}$ and $\mathbf{D}_{x}$ matrices with new values corresponding to the altered topology to calculate voltages. Similarly, in our optimization-based algorithm, adjustments can be made by excluding optimized parameters for removed lines and incorporating original resistance and reactance values for new lines in $\mathbf{D}^{opt}_{r}$ and $\mathbf{D}^{opt}_{x}$, without altering the bias parameters ($\boldsymbol{\gamma}^{opt}$, $\boldsymbol{\rho}^{opt}$, and $\boldsymbol{\varrho}^{opt}$). Successfully maintaining performance with this strategy indicates that our algorithm adapts well to different network topologies, avoiding overfitting to a specific configuration. Alternatively, optimizing parameters specifically for each topology through dedicated optimization could enhance accuracy at the cost of increased computational time and storage for the additional parameters needed. Nevertheless, such calculations could be efficiently executed in parallel for a specified set of topologies, making this process suitable for high-performance computing environments, as each topology’s optimization process operates independently from others.
@@
We next evaluate the adaptability of optimized parameters, i.e., assess how well parameters optimized for one topology perform in others. To accomplish this, we performed the offline training phase of our proposed algorithm using the same setup as before across the $35$ network topologies to obtain a dataset with $35$ sets of optimized coefficients and bias parameters. We tested the performance of these optimized parameters using the same $10,000$ test samples as before on each topology.
@@
The matrix in Fig. 10 further illustrates the performance of our optimized parameters relative to the traditional LinDistFlow approximation (i.e., comparing each row from heatmap in Fig. 9 to the horizontal vector plot at the bottom of the figure). This matrix employs a binary color coding—green for topologies where the optimized parameters outperform the traditional LinDistFlow approximation and white where they do not. For example, row $1$ in this matrix shows that training the parameters using the base topology results in the optimized parameters outperforming the traditional LinDistFlow approximation on topologies $1$, $5$, $6$, $9$, $19$, $21$, $24$, $26$, $29$, and $31$. As another example, training the parameters using topology $2$ leads to superior performance on topologies $\{[2,4]\cup 8\cup[10,15]\cup[17,18]\cup 20\cup[22,23]\cup 25\cup 28\cup 30\cup 33\cup 35\}$.
@@ V-G Illustrative Application: Hosting Capacity Computation
$\min_{p_{n}^{g},q_{n}^{g}}\quad$ $\sum_{n\in\mathcal{\hat{N}}}\left(\frac{(\overline{p_{n}^{g}}-p_{n}^{g})^{2}}{\overline{p_{n}^{g}}}+\xi\frac{{q_{n}^{g}}^{2}}{\overline{s_{n}^{g}}}\right)$ (15a)
s.t. $0\leq p_{n}^{g}\leq\overline{p_{n}^{g}},\quad|q_{n}^{g}|\leq\sqrt{\overline{{s_{n}^{g}}^{2}}-{p_{n}^{g}}^{2}},~{}\forall n\in\mathcal{\hat{N}}$ (15b)
$v_{0}=1,\quad\underline{v_{n}}\leq v_{n}\leq\overline{v_{n}},\hskip 51.21504pt\forall n\in\mathcal{N}^{\prime}$ (15c)
$\sqrt{P^{2}_{n}+Q^{2}_{n}}\leq S_{n},\hskip 41.25641pt\forall n:(\pi_{n},n)\in\mathcal{E}$ (15d)
$\sqrt{P_{T}^{2}+Q_{T}^{2}}\leq S_{T},$ (15e)
$P_{T}=\sum_{n:(0,n)\in\mathcal{E}}P_{0n},\quad Q_{T}=\sum_{n:(0,n)\in\mathcal{E}}Q_{0n},$ (15f)
@@
Upon solving (15) with both LDF and OLDF models, we obtain the optimal active and reactive power settings for the inverter-based generation units. By assessing these optimal settings using the original nonlinear DistFlow model, we compare the performance of the OLDF and LDF approximations. Fig. 11 shows that while the traditional LinDistFlow approximation leads to voltage violations at certain buses within the hosting capacity problem (15), the application of the proposed OLDF model avoids such violations. The voltage violations resulting from the LDF-based optimization in Fig. 11 occur because traditional LDF inaccurately models the system’s voltage response to the high generation injections inherent in hosting capacity analysis. This can lead LDF to identify generation setpoints that appear feasible within its simplified model but are, in fact, infeasible in the true nonlinear system. Conversely, OLDF’s parameters, having been optimized for broad accuracy across diverse conditions, provide a more accurate system representation, thus yielding genuinely feasible solutions. Such discrepancies, where LDF-based decisions prove unreliable when operating near system limits, are generally anticipated given OLDF’s consistently superior performance demonstrated across various test cases and loading scenarios (Tables I, II, and III).
@@ VI Conclusion
Since our optimized formulation has the same linear form as the traditional LinDistFlow approximation and performs well for varying topologies, the accuracy advantages of the proposed approach can be directly exploited in a wide range of other applications such as those in [11, 12, 13, 21, 22, 15, 16, 23, 24, 25]. Our future work aims to explore such applications. Our future work will also focus on extending the OLDF framework to effectively address systems with active control devices, such as voltage regulators and switchable capacitor banks. While the current OLDF enhances the modeling of the passive network (with capacitor bank actions represented by their reactive power injections, $q_{n}$, and fixed regulator effects implicitly learned during training), its optimal integration with dynamically operating controls warrants dedicated investigation. Key research directions include tailoring OLDF training strategies to capture the diverse system states induced by these controls and developing methods for adapting OLDF parameters based on their operational status, potentially drawing insights from our topology adaptation analysis. Such advancements aim to further enhance OLDF’s practical utility in actively controlled distribution systems.

### v4

@@ V-A Algorithm Training
As in [19], we use $20$ power injection scenarios during our algorithm’s training phase. This selection is supported by a sensitivity analysis (see Fig. 3), which showed diminishing returns in accuracy improvement with a larger number of scenarios, yielding a balance between model fidelity and training efficiency. These scenarios scale the nominal power injections at each bus by normally distributed multipliers with a mean of one and standard deviation of $35\%$. For the nonlinear DistFlow solutions, we use PowerModels.jl [54] and OpenDSS with the OpenDSSDirect.py package for single- and three-phase networks, respectively, on a PACE computing node at Georgia Tech with a 24-core CPU and 32 GB of RAM. The training algorithm is implemented in Python 3 in a Jupyter Notebook using the TNC method from scipy.optimize.minimize with objective function (5), Jacobian $\mathbf{g}=[{\mathbf{g}^{D_{r}}}^{\top},{\mathbf{g}^{D_{x}}}^{\top},{\mathbf{g}^{\gamma}}^{\top},{\mathbf{g}^{\rho}}^{\top},{\mathbf{g}^{\varrho}}^{\top}]^{\top}$, convergence tolerance of $1\times 10^{-6}$ per unit, and iteration limit of 100100.222Code is available at https://github.com/BabakTaheri1/OLDF
@@ V-B Performance Metrics
We quantify approximation accuracy by comparing the voltage magnitude outputs (denoted $\mathbf{v}^{[\text{model}]}$, where $\text{model}]$ is OLDF, PLPF, LoDF, LDF, or DLPF for the various approximations) against the nonlinear DistFlow solutions (denoted $\mathbf{v}^{DF}$). We use max and mean error metrics in per unit (p.u.):
$\varepsilon_{\text{max}}^{[\text{model}]}$ $=\max_{m\in\mathcal{M}}\|\mathbf{v}_{m}^{[\text{model}]}-\mathbf{v}_{m}^{\text{DF}}\|_{\infty}$ (13)
$\varepsilon_{\text{avg}}^{[\text{model}]}$ $=\frac{1}{|\mathcal{M}||\mathcal{N}^{\prime}|}\sum_{m\in\mathcal{M}}\|\mathbf{v}_{m}^{[\text{model}]}-\mathbf{v}_{m}^{\text{DF}}\|_{1}$ (14)
@@
Figure 7: Voltage profiles for the base case using the nonlinear DistFlow (OpenDSS in black), LinDistFlow with initial parameters (LDF in green), and proposed OLDF (LinDistFlow with optimized parameters in red).
@@ V-D2 High Load Evaluation
Following the methodology described in [19], we generated high-load scenarios by scaling the base loads with a factor that ranges from $2,-1]\cup[1,2]$ at a granularity of $\frac{1}{14}$, yielding $30$ distinct test scenarios. As shown in Table II, our proposed OLDF algorithm surpasses the others in reducing the average voltage estimation error ($\varepsilon_{avg}$) across nearly all scenarios with the exception of the 22-bus case. For this case, our OLDF results were better than all except the PLPF approximation where the average error was still quite close ($0.00020$ p.u. for PLPF vs. $0.00028$ p.u. for OLDF). Regarding the maximum errors in the high-load scenarios, no individual approximation consistently dominated the others across all test cases. However, we note that summing the maximum errors across all test cases reveals that the OLDF parameters lead to the best performance overall for this metric. These results show that OLDF parameters trained with scenarios around base-load conditions nevertheless perform well for high-load conditions. Furthermore, as an adaptive power flow, the OLDF approximation has the ability to tailor the parameters to perform even better for these conditions by including more training scenarios associated with high loading.
@@ V-E Computational Efficiency
The computation times for online uses of the optimized LinDistFlow parameters depend on the particular application for which they are employed. However, since the only changes are to the parameter values and not the mathematical form of the LinDistFlow expressions, online computation times with our optimized parameters should be comparable to existing LinDistFlow approximations. To show this, the average calculation times for the $10,000$ random load scenarios in the experiment in Section V-D3 ranged from $0.0004$ to $0.1156$ seconds, with differences of less than 22% between the LinDistFlow using traditional parameters and our optimized parameters.
@@ V-F Topology Analysis
Addressing topology changes with the LinDistFlow approximation can be approached in various ways. For example, the traditional LinDistFlow approximation allows for straightforward topology adjustments by updating the $\mathbf{D}_{r}$ and $\mathbf{D}_{x}$ matrices with new values corresponding to the altered topology to calculate voltages. Similarly, in our optimization-based algorithm, adjustments can be made by excluding optimized parameters for removed lines and incorporating original resistance and reactance values for new lines in $\mathbf{D}_{r}^{opt}$ and $\mathbf{D}^{opt}_{x}$, without altering the bias parameters ($\boldsymbol{\gamma}^{opt}$, $\boldsymbol{\rho}^{opt}$, and $\boldsymbol{\varrho}^{opt}$). Successfully maintaining performance with this strategy indicates that our algorithm adapts well to different network topologies, avoiding overfitting to a specific configuration. Alternatively, optimizing parameters specifically for each topology through dedicated optimization could enhance accuracy at the cost of increased computational time and storage for the additional parameters needed. Nevertheless, such calculations could be efficiently executed in parallel for a specified set of topologies, making this process suitable for high-performance computing environments, as each topology’s optimization process operates independently of others.
@@
We next evaluate the adaptability of optimized parameters, i.e., assess how well parameters optimized for one topology perform on others. To accomplish this, we performed the offline training phase of our proposed algorithm using the same setup as before across the $35$ network topologies to obtain a dataset with $35$ sets of optimized coefficients and bias parameters. We tested the performance of these optimized parameters using the same $10,000$ test samples as before on each topology.
@@
The matrix in Fig. 10 further illustrates the performance of our optimized parameters relative to the traditional LinDistFlow approximation (i.e., comparing each row from heatmap in Fig. 9 to the horizontal vector plot at the bottom of the figure). This matrix employs a binary color coding—green for topologies where the optimized parameters outperform the traditional LinDistFlow approximation and white where they do not. For example, row $1$ in this matrix shows that training the parameters using the base topology results in the optimized parameters outperforming the traditional LinDistFlow approximation on topologies $1$, 55, 66, 99, 1919, 2121, $24$, $26$, $29$, and $31$. As another example, training the parameters using topology $2$ leads to superior performance on topologies $\{[2,4]\cup 8\cup[10,15]\cup[17,18]\cup 20\cup[22,23]\cup 25\cup 28\cup 30\cup 33\cup 35\}$.
@@
The adaptability of the OLDF framework also extends to modeling active control devices like switched capacitors and voltage regulators. In scenarios where these devices are in a fixed configuration or are locally controlled, the offline training process has an advantage over traditional methods by implicitly learning the average impact of their behavior from the power flow solution samples. For optimization problems where device settings are decision variables, the OLDF model is also well-suited. The reactive power injections from capacitor banks can be treated as direct inputs ($q_{n}$) to the model. Furthermore, different tap positions of a voltage regulator can be modeled as distinct network configurations, allowing for tailored OLDF parameters to be trained offline for each setting, leveraging the same parallelizable approach used for topology changes. This flexibility allows the OLDF model to create more accurate, system-specific linearizations for a wider range of practical operating scenarios.
@@ V-G Illustrative Application: Hosting Capacity Computation
$\operatorname*{minimize}_{p_{n}^{g},q_{n}^{g}}\quad\sum_{n\in\mathcal{\hat{N}}}\left(\frac{(\overline{p_{n}^{g}}-p_{n}^{g})^{2}}{\overline{p_{n}^{g}}}+\xi\frac{{q_{n}^{g}}^{2}}{\overline{s_{n}^{g}}}\right)$ (15a)
$\text{s.t.}\quad 0\leq p_{n}^{g}\leq\overline{p_{n}^{g}},\quad\quad|q_{n}^{g}|\leq\sqrt{\overline{{s_{n}^{g}}^{2}}-{p_{n}^{g}}^{2}},~\forall n\in\mathcal{\hat{N}}$ (15b)
$\quad\quad v_{0}=1,\quad\underline{v_{n}}\leq v_{n}\leq\overline{v_{n}},\hskip 51.21504pt\forall n\in\mathcal{N}^{\prime}$ (15c)
$\quad\quad\sqrt{P^{2}_{n}+Q^{2}_{n}}\leq S_{n},\hskip 41.25641pt\forall n:(\pi_{n},n)\in\mathcal{E}$ (15d)
$\quad\quad\sqrt{P_{T}^{2}+Q_{T}^{2}}\leq S_{T},$ (15e)
$\quad\quad P_{T}=\sum_{n:(0,n)\in\mathcal{E}}P_{0n},\quad Q_{T}=\sum_{n:(0,n)\in\mathcal{E}}Q_{0n},$ (15f)
@@
Upon solving (15) with both LDF and our proposed OLDF models, we obtain the optimal active and reactive power settings for the inverter-based generation units. By assessing these optimal settings using the original nonlinear DistFlow model, we compare the performance of the OLDF and LDF approximations. To clarify the comparison methodology, we first solved the optimization problem (15) using the LDF model to find its proposed optimal generator setpoints. We then solved the same problem using the OLDF model to find a different set of optimal setpoints. Finally, to determine the true feasibility and performance of each solution, we evaluated both sets of setpoints using the accurate nonlinear DistFlow model. Fig. 11 plots the actual voltage profiles resulting from this validation. As shown in this figure, while the traditional LinDistFlow approximation leads to voltage violations at certain buses within the hosting capacity problem (15), the application of the proposed OLDF model avoids such violations. The voltage violations resulting from the LDF-based optimization in Fig. 11 occur because traditional LDF inaccurately models the system’s voltage response to the high generation injections inherent in hosting capacity analysis. This can lead LDF to identify generation setpoints that appear feasible within its simplified model but are, in fact, infeasible in the true nonlinear system. Conversely, OLDF’s parameters, having been optimized for broad accuracy across diverse conditions, provide a more accurate system representation, thus yielding genuinely feasible solutions. Such discrepancies, where LDF-based decisions prove unreliable when operating near system limits, are generally anticipated given OLDF’s consistently superior performance demonstrated across various test cases and loading scenarios (Tables I, II, and III).
@@ VI Conclusion
Since our optimized formulation has the same linear form as the traditional LinDistFlow approximation and performs well for varying topologies, the accuracy advantages of the proposed approach can be directly exploited in a wide range of other applications such as those in [11, 12, 13, 21, 22, 15, 16, 23, 24, 25]. Our future work aims to explore such applications. Building on prior literature such as [57, 58, 59], our future work will also focus on extending the OLDF framework to more effectively optimize setpoints for active control devices, such as voltage regulators and switchable capacitor banks. While the current OLDF enhances the modeling of the passive network (with capacitor bank actions represented by their reactive power injections, $q_{n}$, and fixed regulator effects implicitly learned during training), its optimal integration with dynamically operating controls warrants dedicated investigation. Key research directions include tailoring OLDF training strategies to capture the diverse system states induced by these controls and developing methods for adapting OLDF parameters based on their operational status, potentially drawing insights from our topology adaptation analysis. Such advancements aim to further enhance OLDF’s practical utility in actively controlled distribution systems.